# X-LeBench: A Benchmark for Extremely Long Egocentric Video Understanding

### v1

@@ 4.1 Settings
To better align with real-world usage, our tasks are defined as requiring free-form textual outputs rather than close-set answers for multiple-choice questions. This approach increases practical usability while significantly raising the complexity of tasks, as it demands more nuanced understanding and reasoning. To ensure the evaluability of outputs, the prompts provide extended context along with task-specific instructions and output format requirements for each task type, standardizing the expected responses. For the temporal localization and summarization tasks, each query is independently evaluated, while for the counting and ordering tasks, we aggregate the query contents of each data entry to action or order lists for a unified assessment. Notably, the computational cost and time associated with analyzing ultra-long contexts, along with their overall poor performance and limited analytical value, we test only a limited set of temporal localization tasks.
Overall, X-LeBench presents significant challenges for current video understanding systems, which often lack robust temporal reasoning or retrieval capabilities and are typically constrained to short contexts or video understanding. In this section, we evaluated X-LeBench using three multimodal approaches:
(1) Gemini-1.5-Flash [39]: This native multimodal model excels in handling ultra-long contexts and is trained jointly on multimodal data. It is suitable to directly evaluate X-LeBench in an end-to-end way. In our experiments, the model is set to a temperature of 0.1, and 1200 frames of each data are uniformly sampled as input context.
(2) Socratic Models [45]: Most advanced multimodal systems cannot process ultra-long videos. Inspired by related works [6, 45], we divide videos into 30-second segments and generate captions for these segments using the Qwen-VL-7B model [41], captioning segments at 1 frame per second and resolution settings of $560\times 420$. The captions are then aggregated with timestamps as life-log records, forming the input for executing long-context understanding tasks. The Gemini-1.5-Flash model is used for the long-context question-answering tasks on these textual life log records, with the same parameters as in method (1).
(3) Retrieve-Socratic: Considering the inherent temporal nature of our dataset, we propose an enhanced version of the Socratic Models approach. This method incorporates a rule-based filtering mechanism that extracts only the time segments relevant to the task, thereby excluding irrelevant temporal information from the life-log records. By narrowing the contextual focus, this approach aims to improve efficiency and relevance in long-context understanding. The extracting contextual information is then processed using the same Socratic Models’ pipeline and parameter settings as in method (2).
Method Data Category Object-related Retrieval (%) People-related Retrieval (%) Moment Retrieval (%) Summarization (RougeL / BertScore) (100) AVG. Temporal Localization (%) AVG. Summarization (100) Action Counting (%) Ordering (%)
single multi-video holistic
Gemini-1.5-Flash Short 1.67 0.00 10.26 25.02 / 88.16 20.44 / 86.81 19.96 / 85.65 4.20 23.47 / 87.53 13.87 93.70
Medium 1.92 6.90 10.53 26.00 / 88.55 17.69 / 85.64 16.10 / 83.37 5.88 24.15 / 87.75 16.74 44.66
Long 1.75 11.11 8.33 26.24 / 88.67 17.46 / 84.83 14.12 / 82.96 6.45 24.72 / 87.98 16.76 24.40
Socratic Models Short 8.33 25.00 17.95 24.12 / 88.21 17.73 / 86.09 22.65 / 87.92 14.28 23.01 / 87.87 9.53 85.59
Medium 5.77 20.69 18.42 24.66 / 88.24 15.66 / 85.34 23.79 / 88.53 13.45 23.51 / 87.92 14.04 55.63
Long 12.28 22.22 10.83 24.53 / 88.13 15.70 / 85.28 21.94 / 87.80 11.83 23.39 / 87.79 19.85 23.33
Retrieve - Socratic Short 8.33 25.00 17.95 23.06 / 88.11 20.60 / 86.78 22.97 / 88.05 14.28 22.74 / 87.93 9.92 87.67
Medium 9.62 17.24 7.89 22.20 / 87.85 18.16 / 86.12 24.18 / 88.67 10.92 21.91 / 87.72 9.38 54.55
Long 10.53 66.67 14.17 22.33 / 87.87 18.36 / 86.28 22.94 / 88.31 15.59 21.98 / 87.74 9.77 24.17
Gemini-1.5-Flash All 1.77 5.17 9.14 25.97 / 88.55 18.28 / 85.60 16.87 / 84.06 5.66 24.30 / 87.82 16.44 41.96
Socratic Models All 8.87 22.41 13.70 24.51 / 88.18 16.07 / 85.45 22.79 / 88.08 12.97 23.36 / 87.86 17.15 42.61
Retrieve - Socratic All 9.47 27.59 13.71 22.39 / 87.90 18.72 / 86.32 23.36 / 88.34 13.92 22.08 / 87.77 9.67 43.01
Table 5: Evaluation results on X-LeBench. Including temporal localization tasks (object-related, people-related and moment retrieval), summarization tasks (single-video, multi-video and holistic level), action counting and summary ordering task. AVG. temporal localization is the average performance of all temporal localization tasks, AVG. summarization is the average performance of all summarization tasks.
Figure 5: Performance comparison on X-LeBench. (a): The performance comparison of each method across tasks for different data categories. (b): The performance comparison across different data categories for different tasks.
4.2 Results
As shown in Fig. 5 and Tab. 5, we report the performance of 3 methods on the “short”, “medium”, “long” and “all” categories of data across different tasks. The “all” category represents an aggregated evaluation of the former 3 data types, offering a more comprehensive and holistic assessment. The evaluation metrics for each task are detailed in Sec. 3.2.
Overall Performance. As shown in last three rows of Tab. 5 and Fig. 5 (a), the Retrieve-Socratic method demonstrates superior performance in most tasks. Notably, in the temporal localization task, the Retrieve-Socratic method achieves an average recall that surpasses Gemini-1.5-Flash by 8.26%. In summarization tasks, Gemini-1.5-Flash exhibits stronger performance, specifically, for the single-video summarization subtask, it outperforms other methods. However, due to input token constraints, longer videos require larger sampling intervals, leading to increasing information loss. As a result, Gemini-1.5-Flash’s performance is notably inferior to that of the Socratic Models and Retrieve-Socratic methods for temporal localization tasks, where temporal information of finer granularity is crucial. In holistic summarization and ordering tasks, the queried time periods are global. Under such conditions, the Socratic Models and Retrieve-Socratic method yield nearly identical results, demonstrating strong consistency in performance across these methods. Interestingly, our analysis of the specific results indicates that the Retrieve-Socratic method, which employs temporal information filtering to reduce contextual information, tends to output 0 more frequently for the counting task. Consequently, its performance on the counting task is notably worse than other methods. This suggests that excessive pruning of contextual information may inadvertently lead to unpredictable result bias.
Impact of Data Types. As illustrated in Fig. 5 (b) and Tab. 5, the performance trends across data types are not uniform. These trends may be influenced by varying task-specific objectives and understanding burden imposed by different data types on each method. Overall, both Socratic and Retrieve-Socratic demonstrate more stable performance trends across data types compared to Gemini-1.5-Flash. This suggests that the processing capability of textual information is more robust when dealing with such long-context information. For the ordering task, all methods achieve strong performance on short videos (accuracy exceeding 85%). However, as video length increases, performance across all methods declines significantly, with accuracy dropping below 25% for long data. For all subtasks, Gemini-1.5-Flash’s performance on long data is worse than on short or medium data, except on people-related retrieval, single-video summarization and counting tasks, Socratic Models and Retrieve-Socratic have the same trends as Gemini-1.5-Flash, except on object-related retrieval and counting tasks, and people- and object-related retrieval tasks, respectively.
Remarks. Notably, due to the information security constraints, Gemini-1.5-Flash refuses to respond to certain data. The refusal rates for short, medium, and long data are 20.14%, 27.08%, and 30.56%, respectively. In contrast, for the Socratic Models and Retrieve-Socratic methods, where video data is first converted into textual information before being processed by Gemini, the refusal rate is reduced to 0%. Additionally, in the multi-video summarization task, a significant portion of responses are invalid even when timestamps are included in input context, and the query time is explicitly specified in the question (e.g., “12:00 to 17:00”). Typical invalid responses include statements like “There is no information available within the specified time period”. We classify such responses as failed replies, which are also excluded from evaluation. The failure rates of the multi-video summarization task are 55.28%, 51.89%, and 57.03% for Gemini-1.5-Flash, Socratic Models and Retrieve-Socratic, respectively.
Summary of Key Findings.
  • Performance on Long Videos: All methods face significant challenges with long video inputs, underscoring the difficulty of handling ultra-long temporal contexts.
  • Retrieve-Socratic Method: Excels in temporal localization tasks, demonstrating strong fine-grained temporal understanding. However, excessive information pruning introduces biases in counting tasks.
  • Gemini-1.5-Flash: Shows superior summarization performance but struggles with temporal localization under long-context inputs due to token limitations and information loss.
  • Performance Across Data Lengths: Performance trends vary across input lengths, with Socratic Models and Retrieve-Socratic exhibiting greater robustness compared to Gemini-1.5-Flash. Suggesting that, under ultra-long video inputs with temporal information, converting information to text is a more efficient and robust way for long-context understanding.
Overall, these results highlight the significant challenges posed by long video input, also emphasize the need for improved cross-length generalization and enhanced temporal information handling.
5 Conclusion
We present a customizable life-logging simulation pipeline for generating realistic and contextually coherent simulated video life logs. Introducing X-LeBench, the first benchmark for ultra-long egocentric video tasks, we curate 432 video life logs from Ego4D, categorized into short, medium, and long durations. X-LeBench also includes a suite of evaluation tasks encompassing temporal localization, summarization, counting, and ordering. By providing this challenging benchmark, we hope to promote the development of advanced models and schemes capable of understanding long-form egocentric videos and tackling the complexities of temporal information.
Despite our efforts to construct an extremely long video dataset with contextually coherent activity contents, limitations remain due to the scarcity and insufficient diversity of available data. For details of videos, contextual inconsistencies exist within the dataset, such as different kitchen or office settings within the same data. However, since our focus is on long-term activity content, these discrepancies are beyond the scope of this work. Future work could involve expanding both the modalities and time spans of collected data through diverse devices and sources to enhance a comprehensive understanding of long-form human activity.

### v2

@@ Appendix C Task Details
C.1 Temporal Localization
Task Design and Annotation: This task evaluates a model’s ability to locate relevant time segments in ultra-long egocentric videos based on user queries. This task, akin to Episodic Memory grauman2022ego4d and Natural Language Video Localization (zhang2021natural; zhang2020span), is uniquely challenging for ultra-long videos due to their extensive duration and diverse content. We divide this task into three subtasks based on the type of information being retrieved: objects, people, and moments. For object- and people-related retrieval, we adopt the annotations from the Episodic Memory task (specifically, the Natural Language Query subtask) in Ego4D. Specifically, we retrieve the annotations corresponding to the videos selected by the daily plan chunks (as generated in Sec. 3.1), classify them into topics based on the query templates defined in Ego4D, and remapped the timestamps to align with the virtual time used in the daily plan.
While Ego4D defines three query categories—objects, people, and place—we argue that the “Place” query template (“Where did I put X?”) fundamentally pertains to object location. Thus, we reclassify these queries under the “object-related” category. Additionally, queries without explicit template annotations were assigned to one of 13 predefined types using an edit-distance-based classification, followed by manual verification. For moment retrieval, annotations for this subtask were derived from the Moments Query subtask of the Episodic Memory task in Ego4D.
Evaluation Metrics: Following previous works on temporal localization grauman2022ego4d; zhang2021natural; zhang2020span, we employed a top-$x$ recall with an intersection over union threshold (“recall@x, tIoU”). $x$ and $t$ are predefined variables that indicate the number of retrieved results we look at and the IoU threshold, respectively. Here, we set $x$ to 5 and $t$ to 0.3.
C.2 Summarization
Task Design and Annotation: Summarization is a fundamental task in numerous benchmarks wang2024lvbench; zhou2024mlvu; ataallah2024infinibench. In our context, it provides insights at varying granularities—specific chunks, predefined periods (e.g., morning, afternoon, evening), and holistic summaries. These diverse scopes challenge systems to summarize at multiple summary levels in ultra-long videos. Our summarization tasks involve generating hierarchical summaries for: Single-video (chunk-level) summaries; Multi-video summaries for predefined periods (morning, afternoon, evening [^1] Morning: before 12 p.m., afternoon: 12 p.m. to 5 p.m., evening: after 5 p.m.); And holistic (full-day) summaries. To generate summaries at these levels, we employed a hierarchical approach inspired by islam2024video. Single-video summaries are initially consolidated during the earlier stages of processing (as described in Sec. 3.1). Then, multi-video summaries are produced by LLM-based aggregation, followed by holistic summaries generated in a similar manner.
Evaluation Metrics: We incorporate LLM-based evaluation metrics inspired by MLVU zhou2024mlvu to evaluate our summary tasks from the perspectives of ’Completeness’ and ’Reliability’.
C.3 Counting
Task Design and Annotation: This task assesses a model’s ability to identify, track, and count occurrences of specific actions across ultra-long videos. The goal is to test fine-grained aggregation capabilities over multiple time segments. Future extensions may include object/person counting.
To maintain annotation reliability, we limit the counting task to intervals with available annotations, as Ego4D annotation only covers a fraction of the total video duration. Unannotated actions could otherwise lead to incorrect counts. Specifically, we derive annotations from the Forecasting task in Ego4D, ensuring consistency.
Table 8: Comparison between our method and random sampling. Left two columns are from our method; Right two columns are from random baseline.
Video Info. Video Summary Timestamp Video Summary Video Info.
Daytime & Indoor
C washes hands and holds
the cupboard door in the bathroom.
06:35
C talks, reads a book, arranges books,
draws, and walks inside a house.
Daytime & Indoor
Daytime & Indoor
C, X, Y, and O discuss
in a meeting room.
08:35
C scrolls through the phone,
climbs a ladder, and fixes cables.
Daytime & Outdoor
Daytime & Indoor
C dials a phone, drinks water,
and writes in a bedroom.
16:35 C plays a drum set in a studio. Nighttime & Indoor
Nighttime & Indoor C reads a book on the bed in the bedroom. 23:00 C prepares avocado in the kitchen. Daytime & Outdoor
C.4 Ordering
Task Design and Annotation: Ordering task evaluates a model’s capability to recognize temporal relationships between main activities across video records in ultra-long video life logs. Unlike prior studies focusing on fine-grained actions, we introduce the novel challenge of ordering events based on single-video summaries. Models are tasked with ordering shuffled single-video summaries of entire life logs, posing significant challenges to the system’s temporal understanding ability in long-context inputs. This setup ensures the initial sequence of events is obscured, requiring the model to reconstruct the correct temporal order.
We formulate the task as an open-ended sequence generation problem. Each video summary receives an index post-randomization, preventing numerical hints about the original order. The model is expected to output the correct order based on its understanding of the contents.
Evaluation Metrics: We use ordering accuracy, defined as the ratio between the number of correctly predicted order and the total number of summaries.
Appendix D Dataset Quality Evaluation
X-LeBench will be made publicly available, users can access and download videos for research under the Ego4D license (see Ego4D website). To validate the quality of our dataset, we conducted a human evaluation focusing on realism and contextual consistency. These key aspects are assessed to ensure the dataset accurately reflects realistic human behavior patterns and maintains logical continuity across activities throughout the day. Evaluators concentrated on the content of the activities, disregarding minor inconsistencies (e.g., clothing or environmental details) that do not affect the overall activity realism. The implementation details of this evaluation are outlined in the following sections.
D.1 Evaluation Criteria
The evaluation criteria are defined as follows:
  • Realism. Each record is assessed for temporal and content alignment with real-world daily activities. Mismatches, such as daytime footage with nighttime timestamps, are considered unrealistic. Evaluators should score all records from 1 (very unrealistic) to 5 (highly realistic).
  • Contextual Consistency. This criterion evaluates the logical flow and smoothness between consecutive records. Abrupt transitions, such as jumping from office work to outdoor activities, are considered a lack of consistency. Evaluators should score all records from 1 (very poor consistency) to 5 (strong sequence). This ensures the dataset reflects realistic daily human behavior.
Figure 6: The content of guideline.
D.2 Evaluation Setting
Given the extremely long duration of the dataset, which makes full evaluation cost-prohibitive, we select 6 data samples (two each from the long, medium, and short categories), consisting of 56 videos, totalling 31.9 hours. The evaluators are provided with a detailed evaluation manual and the original video data. Each evaluator independently assesses the video records and records their results on a formatted debriefing sheet. Kendall’s W abdi2007kendall is computed to assess the inter-rater reliability (IRR) of the evaluations.
D.3 Evaluation Procedure
Evaluator Training. Evaluators are trained using detailed guideline (Fig. 6) that include evaluation standards, examples, and formatting instructions.
Evaluation. Three independent evaluators assess all provided data following the guidance. Evaluators score all records from 1 to 5 for realism and contextual consistency, with a brief justification for records scored below 3.
Consensus Evaluation. Compute inter-rater reliability (IRR) using Kendall’s W for the scores of 56 records to ensure reliability between evaluators. Resolve discrepancies through discussion and re-evaluation. Here, the evaluations achieve 0.536, which suggests a moderate level of agreement among the raters.
Reporting. The final report presents the mean score for all evaluators.
Appendix E Additional Results
E.1 Comparison with Trivial Baseline
Our proposed pipeline is designed to maintain temporal coherence and contextual consistency by integrating multi-dimensional information (time, scene, and content) via life-logging simulation pipeline. To validate the effectiveness of our method, we compare it with a trivial baseline that randomly samples videos. Table 8 shows the comparison between the two methods.
As can be observed, our approach retrieves videos with contextually appropriate timestamps and consistent content, closely mirroring real-life life-logging scenarios. In contrast, random sampling results in time-inconsistent and semantically disjointed videos, further validating the effectiveness of our method.
E.2 Evaluation on LongVU
To extend our evaluation, we tested LongVU shen2024longvu, a state-of-the-art open-source model tailored for long video understanding. Results are summarized in Table 9.
Table 9: Performance of LongVU.
Data Category Holistic Summary (10) Ordering (%)
Short 5.21 37.5
Medium 4.81 1.70
Long 4.73 0.69
Our manual review of outputs shows LongVU struggles with temporal reasoning. Even with explicit cues (e.g., “frames 0–500 are recorded from 13:00–13:10”), it still outputs summaries covering the full video. In ordering tasks, LongVU frequently outputs invalid sequences such as “0, 1, 2, 3” or “100, 100, 100, 100”, especially for long videos. Consequently, we only evaluated LongVU on tasks that require global information or coarse-grained temporal understanding (i.e., holistic summarization and ordering). Its performance significantly drops as video length increases, underscoring the limitations of current open-source models in ultra-long video understanding and temporal reasoning.